OpenStudy (anonymous):

A wooden pyramid, 3 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 6 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

OpenStudy (anonymous):

yeh... its talking yoshi/rayman guy!

OpenStudy (anonymous):

welcome to mcdonalds, please may i take your order?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I want a hot fudge sundae1

OpenStudy (anonymous):

that will be all

OpenStudy (anonymous):

formula: volume of pyramid = l*w*h/3

OpenStudy (anonymous):

k

OpenStudy (anonymous):

l*w = base (square) h = 3

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the dimensions are increased by a factor of 6

OpenStudy (anonymous):

if it is magnified by 6, 3*6 =18

OpenStudy (anonymous):

In order to answer this question, you must recognize the relation between length and volume. The units of volume are (length)^3, thus, in order to find the ratio of volumes, you simply take the factor change of length, 6, and cube it. Thus, the ratio of volumes will be 36.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ehhh...

OpenStudy (anonymous):

uj?

OpenStudy (anonymous):

affirmative

OpenStudy (anonymous):

jelliot is correct

OpenStudy (anonymous):

ehh...

OpenStudy (anonymous):

could you explain it in simpler terms?

OpenStudy (anonymous):

let's see...

OpenStudy (anonymous):

k

OpenStudy (anonymous):

for first volume...